OpenStudy (anonymous):

The length of a rectangle is six times its width. If the area of the rectangle is 150m^2 , find its perimeter.?

OpenStudy (anonymous):

 Let x = the width. Then, that means the length = 6x. Opposite sides of a rectangle are equal, so we can draw the picture above. Also, the area of a rectangle is length * width, so the area of this rectangle is 6x². We're given the area is 150m², so: 6x² = 150m² x² = 25m² x = 5m Perimeter = sum of all sides = x + x + 6x + 6x = 14x Perimeter = 14 * 5m = 70m